OpenStudy (anonymous):

the total horizontal force exerted between the tires of a 1500-kg automobile and the ground is 980 N. if the car starts from rest, how far will it go in 5.0s?

OpenStudy (anonymous):

Simple, do F=ma 1500(KG)*(9.81m/s^2)=14,715 N Now. do sum of all force is F=14,715-980=13,735 Now, look for acceleration at that rate. So F=ma again :) 13,735=(1500)a a=9.15 m/^2 v=v0+at <-- Plug in numbers v= 0+(9.15)(5) =45.75 m/s

OpenStudy (stormfire1):

I think he's right except that the final question was how far would the car go from rest: Xf = Xi + Vi(t) + 1/2 at^2 Xf = 0m + 0m/s + .5(9.15m/s^2)(5^2) = ~114.4m

OpenStudy (anonymous):

Lol I guess I was to busy solving it stormfire has the correct answer since you are calculating the distance and not the velocity at 5 seconds.